# FedP2EFT: Federated Learning to Personalize Parameter Efficient Fine-Tuning for Multilingual LLMs

### v1

@@ 4.2 Results on Text Classification
Apart from Standard FL, we show that FedP2EFT can be plugged into existing personalized FL works that trains both a subset of the pretrained model and personalized layers for each client. Tables 3 and 4 show that FedP2EFT outperforms baselines in almost all cases in our XNLI setup given a base model trained using FedDPA-T (Yang et al., 2024b) and DEPT(SPEC) (Iacob et al., 2025) respectively. In short, FedP2EFT can be integrated into a larger family of existing personalized FL approaches, listed in Section 2, to further improve personalization performance.
$\mathbf{r}$ Approach te ar es en fr zh pt Wins
2 LoRA 0.2354/0.1383/0.1344 0.3364/0.0659/0.0656 0.3871/0.4142/0.3855 0.3345/0.3793/0.3102 0.2884/0.3569/0.2968 0.1078/0.1208/0.1194 0.3835/0.4478/0.4091 0
AdaLoRA 0.2373/0.1428/0.1391 0.3440/0.0668/0.0665 0.3944/0.4273/0.3994 0.3536/0.4042/0.3334 0.2858/0.3528/0.2937 0.1078/0.1226/0.1200 0.3834/0.4514/0.4108 2
BayesTune-LoRA 0.2406/0.1440/0.1410 0.3240/0.0579/0.0576 0.3797/0.4065/0.3781 0.2922/0.2841/0.2302 0.2883/0.3535/0.2927 0.0946/0.1132/0.1119 0.3674/0.4327/0.3932 1
FedL2P 0.2291/0.1356/0.1322 0.3329/0.0687/0.0675 0.3783/0.4034/0.3762 0.3250/0.3667/0.3032 0.2944/0.3614/0.3004 0.0869/0.1173/0.1162 0.3776/0.4439/0.4047 0
FedP2EFT 0.2434/0.1440/0.1403 0.3663/0.0785/0.0764 0.3941/0.4224/0.3928 0.3746/0.4321/0.3610 0.3442/0.4057/0.3318 0.1144/0.1171/0.1161 0.3987/0.4646/0.4201 4
4 LoRA 0.2339/0.1317/0.1282 0.3497/0.0679/0.0667 0.4016/0.4369/0.4077 0.3458/0.3993/0.3282 0.2988/0.3698/0.3044 0.1139/0.1206/0.1183 0.3955/0.4588/0.4186 1
AdaLoRA 0.2331/0.1404/0.1365 0.3350/0.0663/0.0659 0.3877/0.4182/0.3900 0.3482/0.3948/0.3252 0.2886/0.3516/0.2930 0.1091/0.1240/0.1215 0.3816/0.4493/0.4091 0
BayesTune-LoRA 0.2369/0.1426/0.1395 0.3324/0.0609/0.0600 0.3846/0.4111/0.3825 0.3121/0.3195/0.2553 0.2900/0.3557/0.2951 0.1104/0.1181/0.1172 0.3707/0.4386/0.4015 0
FedL2P 0.2298/0.1364/0.1327 0.3376/0.0711/0.0692 0.3797/0.4130/0.3836 0.3392/0.3787/0.3130 0.2938/0.3648/0.3053 0.0974/0.1264/0.1240 0.3876/0.4561/0.4159 1
FedP2EFT 0.2455/0.1495/0.1455 0.3671/0.0749/0.0736 0.4021/0.4333/0.3994 0.3831/0.4400/0.3648 0.3381/0.4004/0.3225 0.1129/0.1212/0.1200 0.4018/0.4618/0.4172 5
8 LoRA 0.2361/0.1368/0.1329 0.3573/0.0708/0.0695 0.4017/0.4341/0.4047 0.3586/0.4182/0.3480 0.3047/0.3667/0.3029 0.1156/0.1260/0.1237 0.3982/0.4605/0.4186 0
AdaLoRA 0.2353/0.1443/0.1399 0.3272/0.0648/0.0645 0.3863/0.4217/0.3922 0.3437/0.3876/0.3183 0.2855/0.3552/0.2929 0.1044/0.1242/0.1216 0.3740/0.4421/0.4038 0
BayesTune-LoRA 0.2397/0.1393/0.1355 0.3444/0.0687/0.0678 0.4031/0.4327/0.4032 0.3294/0.3521/0.2812 0.2962/0.3585/0.3008 0.1130/0.1213/0.1193 0.3844/0.4480/0.4084 0
FedL2P 0.2324/0.1352/0.1316 0.3446/0.0698/0.0681 0.3819/0.4153/0.3855 0.3547/0.4082/0.3362 0.3030/0.3700/0.3076 0.0988/0.1217/0.1199 0.3940/0.4611/0.4201 1
FedP2EFT 0.2431/0.1479/0.1442 0.3713/0.0792/0.0779 0.4077/0.4409/0.4063 0.3844/0.4441/0.3687 0.3440/0.4031/0.3222 0.1156/0.1246/0.1240 0.4009/0.4567/0.4119 6
16 LoRA 0.2413/0.1387/0.1355 0.3605/0.0711/0.0699 0.3864/0.4227/0.3897 0.3603/0.4248/0.3533 0.3275/0.3894/0.3178 0.1194/0.1241/0.1227 0.4025/0.4659/0.4225 1
AdaLoRA 0.2349/0.1388/0.1348 0.3248/0.0659/0.0655 0.3805/0.4141/0.3861 0.3346/0.3702/0.3039 0.2818/0.3554/0.2973 0.1022/0.1207/0.1181 0.3686/0.4379/0.4012 0
BayesTune-LoRA 0.2374/0.1351/0.1310 0.3556/0.0813/0.0795 0.3985/0.4317/0.4013 0.3477/0.3998/0.3295 0.2995/0.3621/0.2965 0.1167/0.1205/0.1186 0.3974/0.4579/0.4153 1
FedL2P 0.2345/0.1417/0.1368 0.3457/0.0643/0.0633 0.3884/0.4185/0.3810 0.3740/0.4420/0.3667 0.3301/0.3752/0.2945 0.0930/0.1223/0.1211 0.3956/0.4543/0.4086 0
FedP2EFT 0.2444/0.1447/0.1406 0.3735/0.0750/0.0740 0.4160/0.4462/0.4105 0.3920/0.4488/0.3725 0.3435/0.3992/0.3204 0.1103/0.1289/0.1270 0.4052/0.4623/0.4177 5
$\mathbf{r}$ Approach te ar es en fr zh pt ru Wins
2 LoRA 0.1553/0.0854/0.0838 0.2425/0.0458/0.0418 0.4275/0.4916/0.4406 0.3248/0.3120/0.2494 0.5513/0.6667/0.6667 0.2489/0.0020/0.0020 0.3610/0.4228/0.4033 0.2242/0.1797/0.1735 0
AdaLoRA 0.1595/0.1108/0.1092 0.2326/0.0721/0.0721 0.4340/0.4954/0.4435 0.3234/0.3102/0.2513 0.5513/0.6667/0.6667 0.2504/0.0020/0.0020 0.3338/0.4121/0.3970 0.2335/0.1758/0.1703 1
BayesTune-LoRA 0.1676/0.0888/0.0858 0.2243/0.0487/0.0457 0.3821/0.4564/0.4089 0.3176/0.3069/0.2437 0.3998/0.3333/0.3333 0.2308/0.0020/0.0020 0.3052/0.3716/0.3553 0.2484/0.1779/0.1773 0
FedL2P 0.1568/0.1156/0.1097 0.2368/0.0496/0.0496 0.4134/0.4810/0.4350 0.3141/0.3050/0.2418 0.3998/0.3333/0.3333 0.2350/0.0020/0.0020 0.3442/0.4012/0.3853 0.2451/0.2010/0.2010 1
FedP2EFT 0.1674/0.0736/0.0720 0.2636/0.0485/0.0485 0.4391/0.5335/0.4796 0.3084/0.3278/0.2718 0.2155/0.2222/0.2222 0.2771/0.0066/0.0066 0.3477/0.4295/0.4110 0.3413/0.2929/0.2662 5
4 LoRA 0.1463/0.0852/0.0758 0.2475/0.0332/0.0332 0.4178/0.4762/0.4269 0.3394/0.3264/0.2648 0.5513/0.6667/0.6667 0.2566/0.0020/0.0020 0.3554/0.425/0.4044 0.2368/0.1553/0.1530 0
AdaLoRA 0.1634/0.0798/0.0782 0.2229/0.0575/0.0575 0.4258/0.4871/0.4387 0.3191/0.3129/0.2519 0.5513/0.6667/0.6667 0.2458/0.0020/0.0020 0.3523/0.4242/0.4067 0.2427/0.1716/0.1654 2
BayesTune-LoRA 0.1689/0.1044/0.1027 0.2308/0.0557/0.0557 0.3911/0.4617/0.4141 0.3219/0.3088/0.2445 0.3998/0.3333/0.3333 0.2349/0.0020/0.0020 0.3233/0.3798/0.3642 0.2437/0.1714/0.1653 1
FedL2P 0.1556/0.0627/0.0611 0.2465/0.0372/0.0331 0.4292/0.4984/0.4445 0.3237/0.3103/0.2498 0.5513/0.6667/0.6667 0.2454/0.0020/0.0020 0.3446/0.4223/0.4056 0.2329/0.1703/0.1625 0
FedP2EFT 0.1618/0.0706/0.0689 0.2621/0.0386/0.0386 0.4619/0.5591/0.5061 0.3194/0.3381/0.2835 0.3998/0.3333/0.3333 0.2898/0.0094/0.0094 0.3508/0.4306/0.4115 0.2538/0.1531/0.1531 4
8 LoRA 0.1600/0.0835/0.0791 0.2669/0.0524/0.0484 0.4312/0.5136/0.4573 0.3352/0.3326/0.2706 0.5664/0.6667/0.6667 0.2564/0.0045/0.0045 0.3465/0.4264/0.4090 0.2275/0.1847/0.1847 1
AdaLoRA 0.1604/0.1045/0.1029 0.2226/0.0606/0.0606 0.4136/0.4889/0.4350 0.3200/0.3082/0.2462 0.3164/0.5000/0.5000 0.2503/0.0020/0.0020 0.3398/0.4203/0.4008 0.2370/0.1687/0.1625 1
BayesTune-LoRA 0.1607/0.0735/0.0688 0.2377/0.0731/0.0731 0.4102/0.4803/0.4352 0.3261/0.3114/0.2458 0.3998/0.3333/0.3333 0.2476/0.0020/0.0020 0.3483/0.3998/0.3817 0.2514/0.1668/0.1612 0
FedL2P 0.1602/0.0586/0.0570 0.2462/0.0447/0.0447 0.4339/0.5007/0.4513 0.3482/0.3281/0.2642 0.3301/0.7460/0.7460 0.2617/0.0020/0.0020 0.3362/0.4262/0.4109 0.2300/0.1542/0.1542 1
FedP2EFT 0.1720/0.0948/0.0890 0.2787/0.0804/0.0804 0.5022/0.5825/0.5168 0.3377/0.3661/0.2984 0.2730/0.2756/0.2756 0.3186/0.0249/0.0241 0.3623/0.4375/0.4158 0.2770/0.1171/0.1171 5
16 LoRA 0.1610/0.0901/0.0885 0.2585/0.0403/0.0371 0.4259/0.5245/0.4700 0.3104/0.3302/0.2775 0.4077/0.3571/0.3571 0.2651/0.0020/0.0020 0.3501/0.4260/0.4065 0.2751/0.2048/0.2048 2
AdaLoRA 0.1412/0.0679/0.0659 0.2119/0.0439/0.0439 0.4149/0.4867/0.4382 0.3241/0.3123/0.2473 0.3764/0.2879/0.2879 0.2421/0.0020/0.0020 0.3308/0.4071/0.3882 0.2260/0.1723/0.1667 0
BayesTune-LoRA 0.1518/0.0641/0.0560 0.2448/0.0318/0.0318 0.4331/0.5021/0.4519 0.3187/0.3128/0.2524 0.3928/0.3148/0.3148 0.2546/0.0020/0.0020 0.3378/0.3921/0.3728 0.2566/0.1867/0.1860 0
FedL2P 0.1613/0.0726/0.0707 0.2605/0.0855/0.0841 0.4181/0.5049/0.4764 0.3383/0.3723/0.3079 0.4065/0.3889/0.3889 0.2547/0.0060/0.0060 0.3262/0.3915/0.3695 0.2574/0.1404/0.1404 2
FedP2EFT 0.1670/0.0826/0.0810 0.2809/0.0751/0.0751 0.4935/0.5715/0.5119 0.3424/0.3800/0.3146 0.2562/0.2626/0.2626 0.3293/0.0364/0.0330 0.3452/0.4280/0.4103 0.3320/0.1856/0.1589 4
4.3 Results on Instruction-Tuning Generation
We evaluate our approach on the more challenging real-world multilingual benchmark, Fed-Aya. Tables 6 and 7 show the average METEOR (Banerjee & Lavie, 2005)/ROUGE-1/ROUGE-L (Lin, 2004) of each language given the off-the-shelf instruction finetuned Llama-3.2-3B (Llama-3.2-3B-Instruct) for seen and unseen clients respectively. Similarly, in Appendix Tables 10 and 11, we show the same tables given a pretrained MobileLLaMA-1.4B model trained using Standard FL with LoRA following the training recipe from FedLLM-Bench (Ye et al., 2024). These two models represent scenarios where the base model may or may not be trained using FL. Similarly to our text classification results, we report “Wins” if an approach has a better performance in at least $2$ out of $3$ metrics.
In all four tables, FedP2EFT outperforms baselines in most scenarios. We also observe that FedL2P underperforms standard baselines in most cases, a phenomenon also observed for our XNLI setup when the base model is trained with FedDPA-T (Tables 3). We hypothesize that the inner-loop optimization in FedL2P fail to reach a stationary point [^4] FedL2P relies on the implicit function theorem for hypergradient computation. due to the inherent task difficulty (Fed-Aya) or a less-performant base model, resulting in a sub-optimal hypergradient and downstream performance.
4.3.1 Limitations of FedP2EFT.
In some cases, FedP2EFT falls short, especially in the recall performance (ROUGE), such as Russian (ru) and French (fr) for unseen clients for both base models. These cases highlight a couple of limitations of our approach: i) ru is not seen by PSG during federated training; there are no ru samples in any clients in the seen pool and ii) none of the clients in the unseen pool have fr as a predominant language (Fig. 3). In the case of ru, there are no other languages that are similar to ru in the seen pool, resulting in worse performance. Hence, we do not expect a similar outcome in datasets with a more diverse pool of clients.
For fr, as the number of predominant language samples are orders of magnitude higher than that of fr samples, the generated $\bm{\lambda}$ are catered towards the predominant language. A simple solution to counteract this limitation is to extend PSG to generate $\bm{\lambda}$ per instance, rather than per client. However, doing so is extremely costly, requiring a forward pass through the PSG for every sample. This calls for novel efficient solutions that can better handle each client’s minority languages and is left as future work.
4.4 Cost of FedP2EFT
Table 5 shows the mean latency, in seconds, and the peak memory usage across 100 runs on the first client in the seen pool for $r=16$ using a single Nvidia A100 GPU. Non-FL baselines do not incur a federated training cost while FL approaches requires training the PSG. Comparing FedL2P and FedP2EFT, FedP2EFT does not require expensive second-order optimization, resulting in better efficiency. We also note that FedL2P needs to be run for every rank while FedP2EFT runs once for all targeted ranks.
During inference, FL-based approaches incur an additional forward pass of base model and the PSG compared to non-FL approaches. Memory-wise, FedP2EFT results in the smallest memory footprint for autoregressive generation as the PSG learns not to attach LoRA modules $\lambda_{l}=0$ on some layers, skipping matmul operations entirely.
@@ 4.4 Cost of FedP2EFT
Figure 4: $\bm{\lambda}$ distance among languages in our XNLI setup. Each block shows the log-scale normalized average Euclidean distances between all pairs of clients’ $\bm{\lambda}$ for two languages (see text). The smaller the distance, the more similar $\bm{\lambda}$ is.
@@ 4.5 Further Analysis
In this section, we further analyze $\bm{\lambda}$ and how they differ across languages. Surprisingly, we find that FedP2EFT learns language-agnostic rank structures. In other words, depending on the task and the base model, the rank structure of $\bm{\lambda}$ is fixed across languages. For instance, in the case where $r=2$, FedP2EFT allocate ranks to dense layers instead of attention blocks. With more budget, e.g.$r=16$, FedP2EFT allocates more rank to either the query attention layer or the value attention layer depending on the setup. We show these rank structures across all setups for $r=2$ and $r=16$ in Appendix Fig. 6-15.
While the rank structure is the same across languages, the rank-wise scales (absolute values of $\bm{\lambda}$) differ. Following FedL2P, we visualize the difference in $\bm{\lambda}$ for different languages using the normalized mean distance, $d(j,k)$, between all clients pairs holding data for languages $j$ and $k$. Fig. 4 and Appendix Fig. 5 show these distances for XNLI and MasakhaNEWS setup respectively. Specifically, the value of each block in each figure is computed as follows: $\log(\frac{d(j,k)}{\sqrt{d(j,j)}\sqrt{d(k,k)}})$. Hence, the smaller the distance, the more similar $\bm{\lambda}$ is between languages. The results are aligned with our intuition that similar languages have similar $\bm{\lambda}$. For instance, the closest language to Urdu (ur) is Arabic (ar), both of which have the closest $\bm{\lambda}$ similarity (Fig. 4); likewise, for Tigrinya (tir) and Amharic (amh) in Appendix Fig. 5. We also observe that unrelated languages have similar $\bm{\lambda}$, e.g., Mandarin (zh) and Vietnamese (vi) share similar $\bm{\lambda}$ with the Indo-European languages (Fig. 4). This finding adds to existing evidence that leveraging dissimilar languages can sometimes benefit particular languages (Ye et al., 2024).
@@ Appendix A Detailed Experimental Setup
For reproducibility and completeness, we provide comprehensive details of all setups, datasets, tasks, models, baselines, and hyperparameters. Code is in the process of being released.
@@ A.1 Tasks, Datasets, and Data Partitioning
XNLI (Conneau et al., 2018) A natural language inference benchmark dataset for evaluating cross-lingual understanding covering 15 diverse languages including both high- and low-resources languages: English, French, Spanish, German, Greek, Bulgarian, Russian, Turkish, Arabic, Vietnamese, Thai, Chinese, Hindi, Swahili and Urdu. XNLI consists of premise-hypothesis pairs, labeled as entailment, contradiction, or neutral across different languages. We sample 2k instances from the XNLI train split and 500 instances from the test split for each pool. The data is then divided equally among 20 clients for each language using the latent Dirichlet allocation (LDA) partition with $\alpha=0.5$. Hence, the total number of clients is 600 (15 languages $\cdot$ 20 clients per language $\cdot$ 2 pools).
MasakhaNEWS (Adelani et al., 2023) A news topic classification benchmark designed to address the lack of resources for African languages. It covers 2 high-resource languages, English and French, and 14 low-resource languages, namely Amharic, Hausa, Igbo, Lingala, Luganda, Naija, Oromo, Rundi, chiShona, Somali, Kiswahili, Tigrinya, isiXhosa, and Yorùbá. Each sample contains a headline, the body text, and one of the 7 labels: business, entertainment, health, politics, religion, sports, or technology. We first combine all samples from the MasakhaNEWS train and validation split to form our train set, and use the MasakhaNEWS test split as our test set. We then split both train and test in each of the 16 languages by half for each pool. Following our XNLI setup, we adopt LDA $\alpha=0.5$ and divide each language’s data equally into 10 clients. Hence, the total number of clients is 320 (16 languages $\cdot$ 10 clients per language $\cdot$ 2 pools). Note that unlike XNLI, the number of samples for each language differs, hence there is quantity skew across clients.
[…]
@@ A.2 Models, Tokenizers, and Data Preprocessing
mBERT (Devlin et al., 2018). We use the pretrained multilingual BERT with its WordPiece tokenizer for all sequence classification experiments, namely all XNLI and MasakhaNEWS setups with various base models. For both datasets, we use a training batch size of 32 and pad input tokens on the right to a max token length of 128 and 256 respectively.
MobileLLaMA-1.4B (Chu et al., 2023). We train a base model with a pretrained MobileLLaMA-1.4B with Standard FL using LoRA in our Fed-Aya setup. We use the default LLaMA tokenizer which is a BPE model based on sentencepiece (Kudo & Richardson, 2018) and adopt the UNK token as the PAD token. During training, we use an effective batch size of $16$ and pad right to the longest token in the batch up to a max token length of 1024. For evaluation, we use a batch size of 8, padding left instead, with greedy sampling up to a max new token length of 1024. We use the Alpaca template to format each prompt:
alpaca_template = """Belowisaninstructionthatdescribesatask.Writearesponsethatappropriatelycompletestherequest.
###Instruction:
###Response:{}{}"""
Llama-3.2-3B (Dubey et al., 2024). We use the off-the-shelf Llama-3.2-3B-Instruct model as our base model and its default tokenizer which is a BPE model based on tiktoken [^5] https://github.com/openai/tiktoken. Training and evaluation hyperparameters are the same as the ones we use for MobileLLaMA. The only two differences are 1) we add a PAD token ‘<pad>’, and 2) we use the Llama 3 instruction template instead:
llama3_instruct_template = """<|begin_of_text|><|start_header_id|>user<|end_header_id|>
{}<|eot_id|><|start_header_id|>assistant<|end_header_id|>

### v2

@@ 4.6 Cost of FedP2EFT
Table 6: Mean latency and memory costs across 100 runs of the first client in the seen pool using an NVIDIA A100 GPU.
@@
Figure 4: Cross-lingual $\bm{\lambda}$ distance in our XNLI setup (See Appendix Fig. 15 for our MasakhaNEWS setup). Each block shows the log-scale normalized average Euclidean distances between all pairs of clients’ $\bm{\lambda}$ in their respective languages. The smaller the distance, the more similar $\bm{\lambda}$ is.
@@ 4.7 Further Analysis
We further analyze $\bm{\lambda}$ and how they differ across languages. Surprisingly, we find that FedP2EFT learns language-agnostic rank structures. In other words, depending on the task and the base model, the rank structure of $\bm{\lambda}$ is fixed across languages. For instance, in the case where $r=2$, FedP2EFT allocate ranks to dense layers instead of attention blocks. With more budget, e.g.r=16r=16, FedP2EFT allocates more rank to either the query attention layer or the value attention layer depending on the setup. We show these rank structures across all setups for $r=2$ and $r=16$ in Appendix Fig. 5-14.
While the rank structure is the same across languages, the rank-wise scales (absolute values of $\bm{\lambda}$) differ. Following FedL2P, we visualize the difference in $\bm{\lambda}$ for different languages using the normalized mean distance, $d(j,k)$, between all clients pairs holding data for languages $j$ and $k$. Fig. 4 and Appendix Fig. 15 show these distances for XNLI and MasakhaNEWS setup respectively. Specifically, the value of each block in each figure is computed as follows: $\log(\frac{d(j,k)}{\sqrt{d(j,j)}\sqrt{d(k,k)}})$. Hence, the smaller the distance, the more similar $\bm{\lambda}$ is between languages. The results are aligned with our intuition that similar languages have similar $\bm{\lambda}$. For instance, the closest language to Urdu (ur) is Arabic (ar), both of which have the closest $\bm{\lambda}$ similarity (Fig. 4); likewise, for Tigrinya (tir) and Amharic (amh) in Appendix Fig. 15. We also observe that unrelated languages have similar $\bm{\lambda}$, e.g., Mandarin (zh) and Vietnamese (vi) share similar $\bm{\lambda}$ with the Indo-European languages (Fig. 4). This finding adds to existing evidence that leveraging dissimilar languages can sometimes benefit particular languages (Ye et al. 2024).
@@ Appendix A Detailed Experimental Setup
For reproducibility and completeness, we provide code, along with comprehensive details of all setups, datasets, tasks, models, baselines, and hyperparameters.
@@ A.1 Tasks, Datasets, and Data Partitioning
XNLI (Conneau et al. 2018). A natural language inference benchmark dataset for evaluating cross-lingual understanding covering 15 diverse languages including both high- and low-resources languages: English, French, Spanish, German, Greek, Bulgarian, Russian, Turkish, Arabic, Vietnamese, Thai, Chinese, Hindi, Swahili and Urdu. XNLI consists of premise-hypothesis pairs, labeled as entailment, contradiction, or neutral across different languages. We sample 2k instances from the XNLI train split and 500 instances from the test split for each pool. The data is then divided equally among 20 clients for each language using the latent Dirichlet allocation (LDA) partition with $\alpha=0.5$. Hence, the total number of clients is 600 (15 languages $\cdot$ 20 clients per language $\cdot$ 2 pools).
MasakhaNEWS (Adelani et al. 2023). A news topic classification benchmark designed to address the lack of resources for African languages. It covers 2 high-resource languages, English and French, and 14 low-resource languages, namely Amharic, Hausa, Igbo, Lingala, Luganda, Naija, Oromo, Rundi, chiShona, Somali, Kiswahili, Tigrinya, isiXhosa, and Yorùbá. Each sample contains a headline, the body text, and one of the 7 labels: business, entertainment, health, politics, religion, sports, or technology. We first combine all samples from the MasakhaNEWS train and validation split to form our train set, and use the MasakhaNEWS test split as our test set. We then split both train and test in each of the 16 languages by half for each pool. Following our XNLI setup, we adopt LDA $\alpha=0.5$ and divide each language’s data equally into 10 clients. Hence, the total number of clients is 320 (16 languages $\cdot$ 10 clients per language $\cdot$ 2 pools). Note that unlike XNLI, the number of samples for each language differs, hence there is quantity skew across clients.
Fed-Aya (Ye et al. 2024). A federated instruction tuning benchmark, based on Aya (Singh et al. 2024), where the data is annotated by contributors and partitioned by annotator ID. Following FedLLM-Bench (Ye et al. 2024), we focus on 6 high-resource languages, English, Spanish, French, Russian, Portuguese, Chinese, and 2 low-resource languages, Arabic and Telugu. Additionally, we filter out the other languages from the dataset. Out of 38 clients, we select 8 for our unseen pool, $\text{client\_ids}=[21,22,23,24,25,26,27,34]$ and the rest goes into our seen pool. Each client can have up to 4 languages where the number of data samples can range from a hundred to over a thousand samples per client.
@@ A.2 Models, Tokenizers, and Data Preprocessing
mBERT (Devlin et al. 2018). We use the pretrained multilingual BERT with its WordPiece tokenizer for all sequence classification experiments, namely all XNLI and MasakhaNEWS setups with various base models. For both datasets, we use a training batch size of 32 and pad input tokens on the right to a max token length of 128 and 256 respectively.
MobileLLaMA-1.4B (Chu et al. 2023). We train a base model with a pretrained MobileLLaMA-1.4B with Standard FL using LoRA in our Fed-Aya setup. We use the default LLaMA tokenizer which is a BPE model based on sentencepiece (Kudo and Richardson 2018) and adopt the UNK token as the PAD token. During training, we use an effective batch size of $16$ and pad right to the longest token in the batch up to a max token length of 1024. For evaluation, we use a batch size of 8, padding left instead, with greedy sampling up to a max new token length of 1024. We use the Alpaca template to format each prompt:
1alpaca_template = """Below is an instruction that describes a task. Write a response that appropriately completes the request.
2
3### Instruction:
4{}
5
6### Response: {}{}"""
Llama-3.2-3B (Dubey et al. 2024). We use the off-the-shelf Llama-3.2-3B-Instruct model as our base model and its default tokenizer which is a BPE model based on tiktoken [^7] https://github.com/openai/tiktoken. Training and evaluation hyperparameters are the same as the ones we use for MobileLLaMA. The only two differences are 1) we add a PAD token ‘<pad>’, and 2) we use the Llama 3 instruction template instead:
1llama3_instruct_template = """<|begin_of_text|><|start_header_id|>user<|end_header_id|>
2
3{}<|eot_id|><|start_header_id|>assistant<|end_header_id|>
4
5{}{}
6"""